Solve The Differential Equation Y''+6y'+9y=16x^2

1
Expert's answer
2022-03-02T11:05:45-0500

Corresponding homogeneous differential equation


"y''+6y'+9y =16x^2"

Characteristic (auxiliary) equation


"r^2+6r+9=0"

"r_1=r_2=-3"

The general solution of the homogeneous differential equation is


"y_h=c_1e^{-3x}+c_2xe^{-3x}"

Find the particular solution of the non homogeneous differential equation


"y_p=Ax^2+Bx+C"

"y_p'=2Ax+B"

"y_p''=2A"

Substitute


"2A+12Ax+6B+9Ax^2+9Bx+9C=16x^2"

"A=16\/9"

"B=-64\/27"

"C=32\/27"

The general solution of the non homogeneous differential equation is


"y=c_1e^{-3x}+c_2xe^{-3x}+\dfrac{16}{9}x^2-\dfrac{64}{27}x+\dfrac{32}{27}"